OpenStudy (anonymous):

To use your calculator to find cot^-1(x)-π/2tan^-1(x).

OpenStudy (dumbcow):

are you graphing this? otherwise you need a x-value

OpenStudy (dumbcow):

cot^-1(x) = tan^-1(1/x) if that helps

OpenStudy (anonymous):

If u want to find "x" then you use cot^-1(x)=1/cot(x) and tan^-1(x)=1/tan(x) cot^-1(x) =Π/2 tan^-1(x) tan(x)=Π/2*[1/tan(x)] tan^2(x)=Π/2 tan(x)=√(Π/2) X=tan^-1[√(Π/2)]